OpenStudy (anonymous):

The perimeter of a basketball court is 102 meters and the length is 6 meters longer than twice the width . What are the length and width?

OpenStudy (anonymous):

Perimeter of a rectangle(basketball court)=2L+2W, where L=Length and W=Width. We are given that L=6(2W)........eq1 and 102=2L+2W......eq2. Substitute eq1 into eq2, then 102=2(12W)+2W. 102=24W+2W 102=26W W=51/13 Finally substitute W=51/13 into eq1 then L=12(51/13)=612/13. Rounded off to 3 decimals L=47.077 and W=3,923